OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies directly as x and y=32 when x=8 The variation constant is k= The equation of variations is y=

OpenStudy (anonymous):

I'm not really sure what you want here. But if i understand correctly, it will be y=4x. thus, when x = 8, it will be y = 4*8 = 32 is that what you want?

OpenStudy (anonymous):

Honestly I'm not really sure either :(, sounds good though thank you!!

OpenStudy (anonymous):

What would the k =?

jimthompson5910 (jim_thompson5910):

y = kx 32=k(8) 32/8 = k k = 4 So the equation is y = 4x

OpenStudy (anonymous):

ok so k=4 and y=kx?

jimthompson5910 (jim_thompson5910):

k=4 and y=4x

jimthompson5910 (jim_thompson5910):

You replace k with 4

OpenStudy (anonymous):

OH ok, I get it now, sorry. Thanks for all your help!! :)